# Tony Lindeberg

### v2

@@ 5.2 Transformation properties of the pure spatial and temporal derivative operators
$\nabla_{(x^{\prime},t^{\prime})}=G^{-T}\,\nabla_{(x,t)}.$ (46)
@@ 5.3.1 Scale-normalized spatial derivatives
To handle the effect of uniform spatial scaling transformations on spatial image data, we can introduce scale-normalized spatial derivatives corresponding to the regular spatial gradient operator $\nabla_{x}=(\partial_{x_{1}},\partial_{x_{2}})^{T}$ according to (Lindeberg 1998 Equation (6)) (here simplified by setting the more general scale normalization parameter to $\gamma=1$)
@@
then it can be shown (Lindeberg 1998 Section 4.1; Lindeberg 2025d Sections 3.1–3.2) that under a uniform spatial scaling transformation of the form (1), for matching values of the spatial scale parameters according to (34)
@@
the corresponding scale-normalized derivatives will be equal at corresponding image points $x^{\prime}=S_{x}\,x$ according to
@@
Here, the first expression (52) represents (regular) scale-normalized gradient operators over the domains after and before the spatial scaling transformation. The second expression (53) represents (regular) scale-normalized Hessian operators $({\cal H}L^{\prime})(x^{\prime};\;s^{\prime})$ and $({\cal H}L)(x;\;s)$ over the domains after and before the image transformation. The third expression (54) represents (regular) scale-normalized directional derivative operators of order $m$ over the domains after and before the geometric transformation.
@@ 5.3.2 Affine-normalized spatial derivatives
To handle the effect of more general spatial affine transformations on spatial image data, one needs to make use of different spatial covariance matrices $\Sigma$ and $\Sigma^{\prime}$ for the spatial receptive fields before and after the spatial affine transformation. For this purpose, three main notions [^5] Notably, in the special case when the spatial covariance matrix $\Sigma$ in the affine Gaussian kernel is equal to a unit matrix $\Sigma=I$, the affine Gaussian kernel in (55) reduces to the isotropic Gaussian matrix in (50), in turn implying that the affine scale-normalized directional derivative operator in (62) reduces to the regular scale-normalized directional derivative operator used in (54). Furthermore, when $\Sigma=I$ the scale-normalized affine gradient operator in (62) similarly reduces to the regular scale-normalized gradient operator used in (52). Similarly, when $\Sigma=I$ the scale-normalized affine Hessian operator in (66) reduces to the regular scale-normalized Hessian operator used in (53). In these respects, the affine-normalized derivative operators in Section 5.3.2 constitute generalizations of the previously used regular (isotropic) scale-normalized derivative operators in Section 5.3.1 from an isotropic spatial scale space representation generated by convolution with rotationally symmetric Gaussian kernels to an anisotropic affine Gaussian scale space generated by convolution with anisotropic affine Gaussian kernels. of affine-normalized spatial derivatives have been proposed in Lindeberg (2025d) Sections 3.3–3.8 for the spatial affine scale-space representation $L\colon{\mathbb{R}}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\rightarrow{\mathbb{R}}$ of any 2-D purely spatial image $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ defined according to
@@
generated by convolution with affine Gaussian kernels $g\colon{\mathbb{R}}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\rightarrow{\mathbb{R}}$ according to (26) having spatial covariance matrices $\Sigma$ not equal to a unit matrix:
@@
thus showing that the set of second-order spatial derivatives before and after the image transformation can up to an undeterminacy with respect to a possibly unknown rotation matrix be perfectly matched, see Figure 11 for an illustration.
@@ 5.3.3 Scale-normalized temporal derivatives
To handle the effect of temporal scaling transformations modelling the effect of temporal or spatio-temporal events occurring either faster or slower relative to a reference view, one can introduce scale-normalized temporal derivatives according to Lindeberg (2017) Equation (6)
@@
where $\tau$ denotes the temporal scale parameter in the temporal scale-space representation $L\colon{\mathbb{R}}\times{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$ of a temporal signal $f\colon{\mathbb{R}}\rightarrow{\mathbb{R}}$ according to
@@
With these definitions (and the more general scale normalization power $\gamma$ in Lindeberg (2017) set to $\gamma=1$), the resulting scale-normalized derivatives do under a temporal scaling transformation of the form (4) become equal at corresponding temporal moments $t^{\prime}=S_{t}\,t$ according to (Lindeberg 2017 Equation (10))
@@ 5.3.4 Scale-normalized velocity-adapted temporal derivatives
then the spatio-temporal scale-space representations $L$ and $L^{\prime}$ before and after the composed geometric image transformation can be perfectly matched according to
$L^{\prime}(x^{\prime},t^{\prime};\;s^{\prime},\tau^{\prime},v^{\prime})=L(x,t;\;s,\tau,v),$ (77)
@@
Figures 1313 illustrate these results in terms of commutative diagrams for spatio-temporal receptive field response under geometric image transformations for the specific spatio-temporal receptive field model
@@ 5.6 Relationships to the influence of illumination variations
Regarding variabilities in image data caused by natural image transformations, we do in this treatment focus on the influence due to geometric image transformations. Regarding the influence of illumination variations, which also constitute a large source to variability in image data, it is, however, interesting to note that according to the theory in Lindeberg (2013) Section 2.3 condensely summarized in Lindeberg (2021b) Section 3.4, it holds that if the image data $f$ used as input to the receptive fields are parameterized in terms of the logarithm of the intensities in the dimension of the incoming energy $\log I(x,y)$ or $\log I(x,y,t)$ according to
@@
Figure 14: Conceptual illustration of how sets of spatial and/or spatio-temporal receptive field responses computed over different values of the filter parameters $(s,\Sigma,v,\tau)$ of the receptive fields can be matched between different views of the same scene or spatio-temporal event under different viewing conditions, by making use of the matching relations in Equations (34)–(37) between the filter parameters for the two image domains before and after the geometric image transformation, under variabilities of the parameters $(S_{x},A,v,S_{t})$ or $(B,v,S_{t})$ of the primitive geometric image transformations in Equations (1)–(4), as combined into composed geometric image transformations according to either Equation (5) or Equation (9) in combination with Equation (6). Such matching of the receptive field responses under geometric image transformations is possible for receptive fields that obey provable covariance properties as exemplified in Section $5.$ (The arrows between the boxes indicate the information flow from the image acquisition stage to the matching stage. Regarding the visualized receptive fields in the middle box, for the spatial receptive fields in the left column, the coordinates are the purely spatial coordinates $(x_{1},x_{2})\in{\mathbb{R}}^{2}$, whereas for the spatio-temporal receptive fields, only one of the spatial coordinates is shown, thus with the spatio-temporal image coordinates $(x_{1},t)\in{\mathbb{R}}\times{\mathbb{R}}$. The receptive fields in the top parts of the middle box are separable over image space or joint space-time and are based on isotropic Gaussian smoothing over the spatial domain with the spatial covariance matrix $\Sigma$ equal to a unit matrix $I$. The purely spatial receptive fields in the left bottom part of the middle box are based on spatial smoothing with non-isotropic affine Gaussian kernels for which the spatial covariance matrix is not equal to a unit matrix. The joint spatio-temporal receptive fields in the right bottom part of the middle box are based on velocity-adapted spatio-temporal smoothing kernels in combination with velocity-adapted temporal derivatives. All the spatio-temporal receptive fields in this figure are based on temporal smoothing with the time-causal limit kernel (22).) (Note also that in an actual implementation of a matching scheme of receptive field responses in this way, one could also conceive that such a matching could be performed based on receptive field responses at higher levels in the visual hierarchy, which would indeed be possible based on the presented theory, if the receptive field responses at the higher levels are computed from the receptive field responses from the responses of the simple cells in a causal feed-forward manner that respects the covariance properties.)
@@
A main result of the above presented theory is that the output from both purely spatial and joint spatio-temporal receptive fields according to generalized versions of the idealized receptive field models (18) and (4.3) can be matched under the composed geometric image transformations according to both the monocular projection model in Equations (81)–(82) and the binocular projection model in Equations (83)–(84). This result holds provided that we allow for sets of parameters $(s,\Sigma,\tau,v)$ and $(s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})$ of the receptive fields $T(s,\Sigma,\tau,v)$ and $T^{\prime}(s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})$ to be varied between the image domains before and after the geometric image transformation, and specifically having the values of the receptive field parameters being matched according to Equations (34)–(37) as functions of the parameters $(S_{t},A,u,S_{t})$ of the composed geometric image transformation.
@@
Since the parameters $(S_{t},A,u,S_{t})$ of the geometric image transformation cannot be expected to be a priori known to a vision system that is to analyze an a priori unknown scene, a general purpose strategy for a vision system could therefore be to expand the receptive fields into a rich set of receptive fields, with the shapes of the receptive fields expanded over the degrees of freedom of the corresponding image transformations. Thereby, it would be possible to match the outputs from populations of receptive field to establish a corresponding matching of the receptive field responses obtained from a particular viewing condition in relation to a learned memory of receptive field responses computed from similar objects and spatio-temporal events under different sets of viewing conditions, see Figure 14 for an illustration.
@@
Given this idealized theory of the relationship between receptive field responses under locally linearized geometric image transformations, one may therefore ask if biological vision has evolved to be able to handle the influence of geometric image transformations on the receptive field responses in a way that is closely related to the results from the presented idealized theory. Specifically, one may ask if the shapes of the receptive fields of simple cells in the primary visual cortex are expanded over the degrees of freedom of (i) uniform spatial scaling transformations, (ii) non-isotropic spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations.
@@ 6.1 Purely spatial variabilities in the shapes of spatial and spatio-temporal receptive fields
In Lindeberg (2025a), this problem is addressed in detail in relation to the first degrees of freedom concerning the combined effect of (i) uniform spatial scaling transformations and (ii) non-isotropic spatial affine transformations. In brief, and extended from a purely spatial domain to also encompass the joint spatio-temporal domain, the results from that treatment are that:
@@
Regarding the degree of freedom corresponding to uniform spatial scaling transformations, the corresponding degree of freedom in terms of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ is special in the sense that the spatial affine Gaussian kernel obeys a semi-group property over spatial scales. Hence, any receptive field response at a coarser spatial scale can be computed by affine Gaussian smoothing of the receptive field responses for any spatial scales. Thereby, a vision system could in principle choose to compute the earliest layers of spatial receptive fields at only the finest spatial scale and nevertheless be able to compute the coarser spatial receptive fields at higher layers in the visual hierarchy. Thus, irrespective of whether the spatial receptive fields corresponding to the simple cells are expanded over the spatial scales, it seems very plausible that the vision system should have the ability to compute visual operations corresponding to spatial scale covariance.
@@
From studies of the orientations selectivity properties of simple cells established from neurophysiological cell recordings by Nauhaus et al. (2008) and Goris et al. (2015), it is clear that the simple and complex cells of monkeys and cat have a substantial variability in orientation selectivity properties, ranging from narrow to wide orientation selectivity properties.
From a theoretical analysis of the orientation selectivity properties of the affine Gaussian derivative and affine Gabor models of visual receptive fields in Lindeberg (2025b), we have established a connection that degree of orientation selectivity increases with the degree of elongation of the receptive fields. For the affine Gaussian derivative and affine Gabor models, that degree of elongation corresponds to the ratio between the eigenvalues of the affine Gaussian kernel used in these idealized models.
In Lindeberg (2025c), we have combined these two sources of biological and theoretical knowledge to propose that these results are consistent with the receptive fields of monkeys and cats spanning a variability in the degree of elongation of the receptive fields. In Lindeberg (2025a), we further show that this degree of freedom of the receptive fields corresponding to the degree of freedom spanned by the ratio between the singular values obtained from a singular value decomposition of the affine transformation matrix $A$.
@@
To span all the 4 degrees of freedom corresponding to the combination of uniform spatial scaling transformations with non-isotropic spatial affine transformations, there is one additional degree of freedom that corresponds to generalizing the idealized receptive field models (18) and (4.3) to not necessarily having the orientation $\varphi$ for computing the directional derivatives $\partial_{\varphi}^{m}$ being parallel to any of the eigendirections of the spatial covariance matrix $A$.

### v4

@@ 5.2 Transformation properties of the pure spatial and temporal derivative operators
$\nabla_{(x^{\prime},t^{\prime})}=G^{-T}\,\nabla_{(x,t)},$ (46)
where
$G=\left(\begin{array}[]{ccc}1&0&-u_{1}\\ 0&1&-u_{2}\\ 0&0&1\end{array}\right)$ (47)
and
$G^{-T}=\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ u_{1}&u_{2}&1\end{array}\right).$ (48)
@@ 5.3.1 Scale-normalized spatial derivatives
To handle the effect of uniform spatial scaling transformations on spatial image data, we can introduce scale-normalized spatial derivatives corresponding to the regular spatial gradient operator $\nabla_{x}=(\partial_{x_{1}},\partial_{x_{2}})^{T}$ according to (Lindeberg 1998 Equation (6)) (here simplified by setting the more general scale normalization parameter $\gamma$ to $\gamma=1$)
@@
then it can be shown (Lindeberg 1998 Section 4.1; Lindeberg 2025d Sections 3.1–3.2) that under a uniform spatial scaling transformation of the form (1)
$f^{\prime}(x^{\prime})=f(x)\quad\quad\mbox{for}\quad\quad x^{\prime}=S_{x}\,x$ (53)
for matching values of the spatial scale parameters according to (34)
@@
the corresponding scale-normalized derivatives for the transformed spatial scale-space representation
$L^{\prime}(\cdot;\;s)=g(\cdot;\;s^{\prime},I)*f^{\prime}(\cdot),$ (55)
will be equal at corresponding image points $x^{\prime}=S_{x}\,x$ according to
@@
Here, the first expression (56) represents (regular) scale-normalized gradient operators over the domains after and before the spatial scaling transformation. The second expression (57) represents (regular) scale-normalized Hessian operators $({\cal H}_{x^{\prime}}L^{\prime})(x^{\prime};\;s^{\prime})$ and $({\cal H}_{x}L)(x;\;s)$ over the domains after and before the image transformation. The third expression (58) represents (regular) scale-normalized directional derivative operators of order $m$ over the domains after and before the geometric transformation.
@@ 5.3.2 Affine-normalized spatial derivatives
To handle the effect of more general spatial affine transformations
$f^{\prime}(x^{\prime})=f(x)\quad\quad\mbox{for}\quad\quad x^{\prime}=A\,x$ (59)
on spatial image data, one needs to make use of different spatial covariance matrices $\Sigma$ and $\Sigma^{\prime}$ for the spatial receptive fields before and after the spatial affine transformation. For this purpose, three main notions [^8] Notably, in the special case when the spatial covariance matrix $\Sigma$ in the affine Gaussian kernel is equal to a unit matrix $\Sigma=I$, the affine Gaussian kernel in (60) reduces to the isotropic Gaussian matrix in (52), in turn implying that the affine scale-normalized directional derivative operator in (70) reduces to the regular scale-normalized directional derivative operator used in (58). Furthermore, when $\Sigma=I$ the scale-normalized affine gradient operator in (70) similarly reduces to the regular scale-normalized gradient operator used in (56). Similarly, when $\Sigma=I$ the scale-normalized affine Hessian operator in (74) reduces to the regular scale-normalized Hessian operator used in (57). In these respects, the affine-normalized derivative operators in Section 5.3.2 constitute generalizations of the previously used regular (isotropic) scale-normalized derivative operators in Section 5.3.1 from an isotropic spatial scale space representation generated by convolution with rotationally symmetric Gaussian kernels to an anisotropic affine Gaussian scale space generated by convolution with anisotropic affine Gaussian kernels. of affine-normalized spatial derivatives have been proposed in Lindeberg (2025d) Sections 3.3–3.8.
The theoretical background to these affine-normalized spatial derivatives is that we consider the spatial affine scale-space representation $L\colon{\mathbb{R}}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\rightarrow{\mathbb{R}}$ of any 2-D purely spatial image $f\colon{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ defined according to
@@
generated by convolution with affine Gaussian kernels $g\colon{\mathbb{R}}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\rightarrow{\mathbb{R}}$ according to (26) and having spatial covariance matrices $\Sigma$ not necessarily equal to a unit matrix, with its corresponding transformed affine spatial scale-space representation
$L^{\prime}(\cdot;\;s^{\prime},\Sigma^{\prime})=g(\cdot;\;s^{\prime},\Sigma^{\prime})*f(\cdot),$ (61)
which obeys the affine covariance property (Lindeberg and Gårding 1997 Equation (29))
$L^{\prime}(x^{\prime};\;s^{\prime},\Sigma^{\prime})=L(x;\;s,\Sigma)$ (62)
for matching values of the scale parameters $s$ and $s^{\prime}$ and the spatial covariance matrices $\Sigma$ and $\Sigma^{\prime}$ according to (Lindeberg 2025d Equation (118))
$s^{\prime}\,\Sigma^{\prime}=s\,(S_{x}\,A)\,\Sigma\,(S_{x}A)^{T}=s\,S_{x}^{2}\,A\,\Sigma\,A^{T}.$ (63)
Then, given these affine Gaussian scale-space representations over the domains before and after the affine transformation, affine-normalized spatial derivative operators can be defined as follows:
@@
thus showing that the set of second-order spatial derivatives before and after the image transformation can up to an indeterminacy with respect to a possibly unknown rotation matrix be perfectly matched, see Figure 11 for an illustration.
@@ 5.3.3 Scale-normalized temporal derivatives
To handle the effect of temporal scaling transformations, that model the effect of temporal or spatio-temporal events occurring either faster or slower relative to a reference view, one can introduce scale-normalized temporal derivatives according to Lindeberg (2017) Equation (6)
@@
where $\tau=\sigma_{t}^{2}$ denotes the temporal scale parameter in the temporal scale-space representation $L\colon{\mathbb{R}}\times{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$ of a temporal signal $f\colon{\mathbb{R}}\rightarrow{\mathbb{R}}$ according to
@@
With these definitions (and with the more general scale normalization power $\gamma$ in Lindeberg (2017) set to $\gamma=1$), the resulting scale-normalized derivatives of the transformed temporal scale-space representation
$L^{\prime}_{{t^{\prime}}^{n},\text{norm}}(t^{\prime};\;\tau^{\prime})={\tau^{\prime}}^{n/2}\,\partial_{t^{n}}(h(\cdot;\,\tau^{\prime})*f^{\prime}(\cdot))(t^{\prime};\;\tau^{\prime})$ (79)
do under a temporal scaling transformation of the form (4)
$f^{\prime}(t^{\prime})=f(t)\quad\quad\mbox{for}\quad\quad t^{\prime}=S_{t}\,t$ (80)
become equal at corresponding temporal moments $t^{\prime}=S_{t}\,t$ according to (Lindeberg 2017 Equation (10))
@@ 5.3.4 Scale-normalized velocity-adapted temporal derivatives
then if we define the spatio-temporal scale-space representations $L$ and $L^{\prime}$ before and after the composed geometric image transformation
$f^{\prime}(x^{\prime},t^{\prime})=f(x,t)$ (87)
according to
$\begin{split}L(\cdot,\cdot;\,s,\tau,v)=T(\cdot,\cdot;\,s,\tau,v)*f(\cdot;\cdot),\end{split}$ (88)
$\begin{split}L^{\prime}(\cdot,\cdot;\,s^{\prime},\tau^{\prime},v^{\prime})=T(\cdot,\cdot;\,s^{\prime},\tau^{\prime},v^{\prime})*f^{\prime}(\cdot;\cdot),\end{split}$ (89)
the corresponding scale-normalized velocity-adapted temporal derivatives can be perfectly matched according to
$L^{\prime}_{{\bar{t^{\prime}}^{n}},\scriptsize\mbox{norm}}(x^{\prime},t^{\prime};\;s^{\prime},\tau^{\prime},v^{\prime})=L_{{\bar{t}}^{n},\scriptsize\mbox{norm}}(x,t;\;s,\tau,v),$ (90)
@@
Figures 1313 illustrate these results in terms of commutative diagrams for spatio-temporal receptive field responses under geometric image transformations for the specific spatio-temporal receptive field model
@@ 5.6 Relationships to the influence of illumination variations
Regarding variabilities in image data caused by natural image transformations, we do in this treatment focus on the influence due to geometric image transformations. Regarding the influence of illumination variations, which also constitute a large source to variability in image data, it is, however, interesting to note that according to the theory in Lindeberg (2013a) Section 2.3 condensely summarized in Lindeberg (2021b) Section 3.4, it holds that if the image data $f$ used as input to the receptive fields are parameterized in terms of the logarithm of the intensities in the dimension of the incoming energy $\log I(x,y)$ or $\log I(x,y,t)$ according to
@@
Figure 14: Conceptual illustration of how sets of spatial and/or spatio-temporal receptive field responses computed over different values of the filter parameters $(s,\Sigma,v,\tau)$ of the receptive fields can be matched between different views of the same scene or spatio-temporal event under different viewing conditions, by making use of the matching relations in Equations (34)–(37) between the filter parameters for the two image domains before and after the geometric image transformation, under variabilities of the parameters $(S_{x},A,v,S_{t})$ or $(B,v,S_{t})$ of the primitive geometric image transformations in Equations (1)–(4), as combined into composed geometric image transformations according to either Equation (5) or Equation (9) in combination with Equation (6). Such matching of the receptive field responses under geometric image transformations is possible for receptive fields that obey provable covariance properties as exemplified in Section $5.$ (The arrows between the boxes indicate the information flow from the image acquisition stage to the matching stage. Regarding the visualized receptive fields in the middle box, for the spatial receptive fields in the left column, the coordinates are the purely spatial coordinates $(x_{1},x_{2})\in{\mathbb{R}}^{2}$, whereas for the spatio-temporal receptive fields, only one of the spatial coordinates is shown, thus with the spatio-temporal image coordinates $(x_{1},t)\in{\mathbb{R}}\times{\mathbb{R}}$. The receptive fields in the top parts of the middle box are separable over image space or joint space-time and are based on isotropic Gaussian smoothing over the spatial domain with the spatial covariance matrix $\Sigma$ equal to a unit matrix $I$. The purely spatial receptive fields in the left bottom part of the middle box are based on spatial smoothing with non-isotropic affine Gaussian kernels for which the spatial covariance matrix is not equal to a unit matrix. The joint spatio-temporal receptive fields in the right bottom part of the middle box are based on velocity-adapted spatio-temporal smoothing kernels in combination with velocity-adapted temporal derivatives. All the spatio-temporal receptive fields in this figure are based on temporal smoothing with the time-causal limit kernel (22).)
@@
A main result of the above presented theory is that the output from both purely spatial and joint spatio-temporal receptive fields according to generalized versions of the idealized receptive field models (18) and (4.3) can be matched under the composed geometric image transformations according to both the monocular projection model in Equations (94)–(95) and the binocular projection model in Equations (96)–(97). This result holds provided that we allow for the sets of the parameters $(s,\Sigma,\tau,v)$ and $(s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})$ of the receptive fields $T(s,\Sigma,\tau,v)$ and $T^{\prime}(s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})$ to be varied between the image domains before and after the geometric image transformation, and specifically having the values of the receptive field parameters being matched according to Equations (34)–(37) as functions of the parameters $(S_{t},A,u,S_{t})$ of the composed geometric image transformation.
@@
Since the parameters $(S_{t},A,u,S_{t})$ of the geometric image transformation cannot be expected to be a priori known to a vision system that is to analyze an a priori unknown scene, a general purpose strategy for a vision system could therefore be to expand the receptive fields over a rich set of receptive fields, with the shapes of the receptive fields expanded over the degrees of freedom of the corresponding image transformations. Thereby, it would be possible to match the outputs from populations of receptive fields, to establish a corresponding matching of the receptive field responses obtained from a particular viewing condition in relation to a learned memory of receptive field responses computed from similar objects and spatio-temporal events under different sets of viewing conditions, see Figure 14 for an illustration. [^10] Note also that in an actual implementation of a matching scheme of receptive field responses of this form, one could also conceive that such a matching could be performed based on receptive field responses at higher levels in the visual hierarchy, which would indeed be possible based on the presented theory, if the receptive field responses at the higher levels are computed from the receptive field responses from the responses of the simple cells in a causal feed-forward manner that respects the covariance properties.
@@
Given this idealized theory, one may therefore ask if biological vision has evolved similar mechanisms to be able to handle the influence of geometric image transformations on the receptive field responses in a way that is closely related to the results from the presented theory. Specifically, one may ask if the shapes of the receptive fields of the simple cells in the primary visual cortex are expanded over the degrees of freedom of (i) uniform spatial scaling transformations, (ii) non-isotropic spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations, and if so for what species?
Given evolutionary arguments, one could argue that for at least higher species with more developed general purpose vision systems for addressing a wide range of visual tasks, like primates and cats, it would constitute an evolutionary advantage to adapt their visual system to the structure of the external environment, and thus to the influence of geometric image transformations on the image data that reach their visual sensors. One may therefore ask if there for such higher species under a sufficient evolutionary pressure would be variabilities in the shapes of their visual receptive fields in the visual area corresponding to V1, as implied by what would be the result of expanding the receptive field shapes with respect to the degrees of freedom of the 4 main classes of primitive geometric image transformations.
@@ 6.1 Purely spatial variabilities in the shapes of spatial and spatio-temporal receptive fields
In Lindeberg (2025a), the theoretical problem regarding variabilities in receptive field shapes as implied by covariance properties under geometric image transformations is addressed in detail in relation to the first degrees of freedom concerning the combined effect of (i) uniform spatial scaling transformations and (ii) non-isotropic spatial affine transformations. In brief, and extended from a purely spatial domain to also encompass the joint spatio-temporal domain, the results from that treatment are that:
@@
Regarding the degree of freedom corresponding to uniform spatial scaling transformations, the corresponding degree of freedom in terms of the spatial scale parameter $\sigma_{x}=\sqrt{s}$ is special in the sense that the spatial affine Gaussian kernel obeys a semi-group property over spatial scales (see Lindeberg (1993) Equation (15.35))
$g(\cdot;\;\Sigma_{1})*g(\cdot;\;\Sigma_{2})=g(\cdot;\;\Sigma_{1}+\Sigma_{2}),$ (128)
from which it follows that any receptive field response at a coarser spatial scale, here as represented by the spatial gradient vector $\nabla_{x}L$ or the Hessian matrix ${\cal H}_{x}L$ for the spatial covariance matrix $\Sigma_{2}$, can be computed by affine Gaussian smoothing of the receptive field responses for any finer spatial scale, here as represented by the spatial gradient vector $\nabla_{x}L$ or the Hessian matrix ${\cal H}_{x}L$ defined for the spatial covariance matrix $\Sigma_{1}$
$\begin{split}(\nabla_{x}\,L)(\cdot;\;\Sigma_{2})=g(\cdot;\;\Sigma_{2}-\Sigma_{1})*(\nabla_{x}\,L)(\cdot;\;\Sigma_{1}),\end{split}$ (129)
$\begin{split}({\cal H}_{x}\,L)(\cdot;\;\Sigma_{2})=g(\cdot;\;\Sigma_{2}-\Sigma_{1})*({\cal H}_{x}\,L)(\cdot;\;\Sigma_{1}),\end{split}$ (130)
provided that the difference between the spatial covariance matrices
$\Delta\Sigma=\Sigma_{2}-\Sigma_{1}$ (131)
is a symmetric positive definite matrix.
Thereby, a vision system could in principle choose to compute the earliest layers of spatial receptive fields at only the finest spatial scale and nevertheless be able to compute the coarser spatial receptive fields at higher layers in the visual hierarchy. Thus, irrespective of whether the spatial receptive fields corresponding to the simple cells are expanded over the spatial scales, it seems very plausible that the vision system should have the ability to compute visual operations corresponding to spatial scale covariance.
@@
Figure 15: Orientation map in the primary visual cortex of cat, as recorded by Koch et al. (2016) (OpenAccess), and and demonstrating that the visual cortex of higher mammals can be regarded as performing an explicit expansion of the receptive field shapes over spatial image orientations, as would be the result of combining the notion of covariance over the subgroup of pure image rotations.
@@
From studies of the orientation selectivity properties of simple cells established from neurophysiological cell recordings by Nauhaus et al. (2008) and Goris et al. (2015), it is clear that the simple and complex cells of primates and cat have a substantial variability in orientation selectivity properties, ranging from narrow to wide orientation selectivity properties, as illustrated in Figure 16.
Figure 16: Schematic depiction of results from measurements of the orientation tuning of neurons, at different positions in the visual cortex, adapted from Nauhaus et al. (2008), showing how the orientation tuning changes from broad to sharp, and thus higher degree of orientation selectivity, with increasing distance from the pinwheels, consistent with the qualitative behaviour that would be obtained if the ratio $\kappa$, between the scale parameters in underlying affine Gaussian smoothing step in the idealized models of spatial and spatio-temporal receptive fields, would increase when moving away from the centers of the pinwheels on the cortical surface. (From Lindeberg (2025c) with permission (OpenAccess).)
First-order simple cells Second-order simple cells Complex cells
Refer to caption Refer to caption Refer to caption
Figure 17: Graphs of the orientation selectivity for the idealized models of (top left) simple cells in terms of first-order directional derivatives of affine Gaussian kernels, (top middle) simple cells in terms of second-order directional derivatives of affine Gaussian kernels, (top right) complex cells in terms of directional quasi-quadrature measures that combine the first- and second-order simple cell responses in a Euclidean way. The parameter $\kappa$ represents the degree of elongation of the receptive fields, and corresponds to the ratio between the square roots of the eigenvalues of the covariance matrix $\Sigma$ of the affine Gaussian derivative-based receptive fields. (Horizontal axes: orientation $\theta\in[-\pi/2,\pi/2]$. Vertical axes: Amplitude of the receptive field response relative to the maximum response obtained for $\theta=0$.) (Adapted from Lindeberg (2025c) with permission (OpenAccess).)
From a theoretical analysis of the orientation selectivity properties of the affine Gaussian derivative and affine Gabor models of visual receptive fields in Lindeberg (2025b), we have established a connection that the degree of orientation selectivity increases with the degree of elongation of the receptive fields, see Figure 17. For the affine Gaussian derivative and affine Gabor models, that degree of elongation corresponds to the ratio between the eigenvalues of the affine Gaussian kernel used in these idealized models.
In Lindeberg (2025c), we have combined these two sources of biological and theoretical knowledge to propose that these results are consistent with the receptive fields of primates and cats spanning a variability in the degree of elongation of the receptive fields. In Lindeberg (2025a), we further show that this degree of freedom of the receptive fields corresponding to the degree of freedom spanned by the ratio between the singular values obtained from a singular value decomposition of the affine transformation matrix $A$.
@@
To span all the 4 degrees of freedom corresponding to the combination of uniform spatial scaling transformations with non-isotropic spatial affine transformations, there is one additional degree of freedom that corresponds to generalizing the idealized receptive field models (18) and (4.3) to not necessarily having the orientation $\varphi$ for computing the directional derivatives $\partial_{\varphi}^{m}$ being parallel to any of the eigendirections of the spatial covariance matrix $\Sigma_{\varphi}$.